Consider the real vector space
A = {(a, b, c, d) I a, b, c, d belongs to R,2a + 3b = c + d}.
Find dim A. Also find two distinct subspaces
B1. and B2 of R⁴ such that
A (direct sum) B1=R⁴=A(direct sum)B2
1
Expert's answer
2020-11-23T08:53:35-0500

Let us choose standard basis of "\mathbb{R}^4" and give the linear map "T" from "\mathbb{R}^4\longrightarrow \mathbb{R}" as "e_1\mapsto 2e_1, e_2\mapsto 3e_2, e_3\mapsto -e_3, e_4\mapsto -e_4." Then clearly, "T(a,b,c,d)=2a+3b-c-d." Then A is the kernel of T. Clearly, "T" being non-zero and image being in "\mathbb{R}" its rank is 1. Hence by rank-nullity the dimension of A is 4-1=3. Now "2\times 1+3\times 0=2\neq 0+0" . Hence "e_1 \notin A." Similarly "e_2\notin A." Let "B_i=\langle e_i\rangle" for "i=1,2." Now "A+B_i" is a subspace of "\mathbb{R}^4" . Also "A\cap B_i=\{0\}." Hence "A+B_{i}=A\oplus B_i" . Dimension of "A+B_i=" dim"A+" dim "B_i-" dim "(A\cap B_i)" =3+1-0=4 . But the only subspace of "\mathbb{R}^4" of dimension 4 is "\mathbb{R}^4" itself. Hence "A\oplus B_1=\mathbb{R}^4=A\oplus B_2." Now "\lambda e_1=(\lambda,0,0,0)\neq (0,\mu,0,0)=\mu e_2." Hence "B_1,B_2" are distinct.